OpenStudy (anonymous):

f'(x)=(x^3-15^3+x^.5)/x^2

zepdrix (zepdrix):

What's up bunch? c: Looking for f(x)?

OpenStudy (anonymous):

looking for derivative

OpenStudy (anonymous):

inbox me in i would help u

zepdrix (zepdrix):

\[\Large\rm f(x)=\frac{x^3-15^3+x^.5}{x^2}\]Derivative of this?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Split up the fractions first,\[\Large\rm f(x)=\frac{x^3}{x^2}-\frac{15^3}{x^2}+\frac{x^{1/2}}{x^2}\]

zepdrix (zepdrix):

Simplify each term,\[\Large\rm f(x)=x-15^3x^{-2}+x^{-3/2}\]Do you understand that step lady bunch? We haven't taken a derivative yet, we're just getting it ready.

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Alternatively, you could apply the quotient rule from the start, if you prefer that.

zepdrix (zepdrix):

So from here, it's not too bad. You apply your power rule to each term.

OpenStudy (anonymous):

can u help me zepdrix